Sides of a right angle triangle r in A.P.,if smallest side has length 2 units,find sumof other lengths?

Grade:11

2 Answers

JAYAVARDHAN BAIRABATHINA
10 years ago

h2 = a2 + b2 is only true for right-angled triangles. If all the angles of a triangle are less than 90° then h2 < a2 + b2 For example, in an equilateral triangle with sides 1 1 1 and all angles 60° 12 = 1 < 12 + 12 = 2 If one the angles of a triangle is greater than 90° then h2 > a2 + b2 Note that in any triangle, the longest side h cannot be longer than the sum of the other two sides. so h < a + b. If it equalled the sum of the other two then the triangle is just a line of length a + b = h! For example, if the two shorter sides of a right-angled triangle are 2 cm and 3 cm, what is the length of the longest side? If the longest side is h, then, by Pythagoras'' Theorem, we have: h2 = 22 + 32 = 13 h = √13 = 3·60555

solve by using this simple concept

vamsi
10 years ago

ANSWER IS 6.

LET ANOTHER TWO SIDES OF RIGHT ANGLED TRIANGLE ARE a,b.

BY THE GIVEN SUM

2,a,b ARE IN A.P.

2<a<b

therefore b is the hypotenus of the given right angled triangle

2a=2+b-------(1)

b=√(4+a^2)-----(2)

(2) in (1)

2a=2+√(4+a^2)

by solving

a=8/3

put a value in (1)

therefore b=10/3

sum of the remaining sides =a+b=6